OpenStudy (anonymous):

Helpp! Convert the polar eqn to rectangular form please:) r^2=8/(2-sin^2(feta))

OpenStudy (amistre64):

new pic .... nice :)

OpenStudy (anonymous):

\[r ^{2}=8\div(2-\sin ^{2}\Theta)\]

OpenStudy (amistre64):

r^2 = x^2 + y^2 is good to keep in mind; whether its useful of not I dunno

OpenStudy (anonymous):

thnks:)

OpenStudy (amistre64):

y = r sin(t) is aother .....

OpenStudy (anonymous):

amistre always the flirt....but i dont think sanhitha u can get right answers if u befuddle us with gorgeous pics:)

OpenStudy (anonymous):

haha..tht wasnt rlly my intention..was gettin tired of lookin at tht lil bird..lol

OpenStudy (anonymous):

(x^2 + y^2)^2 = 8/ (2 - y^2/(x^2+y^2))

OpenStudy (amistre64):

lol .... the testosterone clouds the reasoning and him goes all primal lol

OpenStudy (anonymous):

o goddd

OpenStudy (anonymous):

stay on topic u guys

OpenStudy (amistre64):

umm... what was the topic again ;)

OpenStudy (anonymous):

Just multiply it over giving r^2(2-sin^2(theta))=8 2r^2-r^2sin^2(theta)=8 2x^2+2y^2-y^2=8 2x^2+y^2=8

OpenStudy (amistre64):

r2=8÷(2−sin2Θ) 2- sin^2 = 8/r^2 .... yeah like that

OpenStudy (anonymous):

Start by multyplying both sides by \(2-\sin^2\theta\), you get \(2r^2-r^2\sin^2\theta=8 \implies 2x^2+2y^2-y^2=8\). So the equation in rectangular form is \(2x^2+y^2=8\).

OpenStudy (anonymous):

ohh..so we basically cant have any trig in our rectangular eqn..n sin^2 is just uhh..squaring the r?..sry im not so good with this stuff

OpenStudy (anonymous):

never mind wat i jus said

OpenStudy (anonymous):

see didnt i tell u

OpenStudy (anonymous):

wat did u tell me?

OpenStudy (anonymous):

so its 2x^2 +y^2=8?

OpenStudy (anonymous):

just replace r^2 by x^2 + y^2

OpenStudy (anonymous):

and sin theta by y/ sqrt (x^2 + y^2)

OpenStudy (anonymous):

but sin^2 theta wud b y^2/x^2+y^2

OpenStudy (anonymous):

\[(x^{2} + y^{2})(2 - y^2/(x^2 + y^2))\]

OpenStudy (anonymous):

= 8

OpenStudy (anonymous):

ur right

OpenStudy (anonymous):

ohhok..gotcha! thanks a lot u all!:)

OpenStudy (anonymous):

btw is that really you or some actress in the pic? :)

OpenStudy (amistre64):

i got this pic on ebay ;)

OpenStudy (anonymous):

hahah...u think it looks lik an actress??

OpenStudy (anonymous):

definitely...one who makes my knees weak even with a minuscule picture that it is ;)

OpenStudy (anonymous):

u did?!?!

OpenStudy (amistre64):

yep; only paid 126.32 cents for it ;)

OpenStudy (amistre64):

i dont recall why I wanted it tho lol

OpenStudy (anonymous):

didnt kno it was on ebay! lolol

OpenStudy (anonymous):

u indian?

OpenStudy (amistre64):

nah, italian

OpenStudy (anonymous):

do i look lik 1?

OpenStudy (anonymous):

ur name suggests

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

where in south india?

OpenStudy (anonymous):

i dont liv der..nri

OpenStudy (anonymous):

oh..thats explain the bonny nature

OpenStudy (anonymous):

bonny??

OpenStudy (anonymous):

means beautiful..old irish term

OpenStudy (anonymous):

ohhhh...nyway..gatta get bac 2 math

OpenStudy (anonymous):

fyn..good luck..dont hesitate to ask fr help :)

OpenStudy (anonymous):

haha i wont! thnks